PKTYPE(1) PKTYPE(1) NAME pktype−verifyandtranslateapackedfontbitmapfiletoplaintext SYNOPSIS pktypepkfile[outputfile] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thepktypeprogramtranslatesapackedfontfile(pk)(outputby,forexample,gftopk(1)toa filethathumanscanread.Italsoservesasapkfile-validatingprogram(i.e.,ifpktypecanread it,it’scorrect)andasanexampleofapk-readingprogramforothersoftwarethatwantstoread pkfiles. Thepkfileonthecommandlinemustbecomplete.Becausetheresolutionispartoftheextension,itwouldnotmakesensetoappendadefaultextensionasisdonewithTEXorDVI-reading software.Ifnooutputfileisspecified,theplaintexttranslationiswrittentostandardoutput. Theoutputfilegivesacompactencodingofthepackedencoding,usingconventionsdescribedin thesourcecode.Runlengthsofblackpixelsalternatewithparenthesizedrunlengthsofwhitepixels,andbracketsareusedtoindicatewhenarowshouldberepeated. ENVIRONMENT pktypelooksforpkfileusingtheenvironmentvariablePKFONTS.Ifthatisnotset,itusesthe variableTEXPKS.Ifthatisnotset,itusesTEXFONTS.Ifthatisnotset,itusesthesystem default. Seetex(1)forthedetailsofthesearching. SEEALSO dvitype(1),gftopk(1),gftype(1). DonaldE.Knuthetal.,METAFONTware. AUTHORS TomasRokickiwrotetheprogram.DonaldE.KnuthoriginallyportedittoUnix.PierreMacKay adapteditforcompilationwithWeb2c. Web2C7.32x 13March1996 1